# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 8 Structural stability analysis
The notion of structural stability of dynamical systems was first introduced by Andronov and Pontryagin under the name coarse (or rough) systems [17]. Different from the previous notions of stability which consider perturbations of initial conditions for a fixed dynamical system, structural stability concerns whether the qualitative behavior of the system trajectories will be affected by small perturbations of the system model itself [41, 191, 232, 250, 279]. In this section, we first discuss the mathematical definition of structural stability in dynamical systems. Then, we present various metrics that can be used to quantify the structural stability of ecological systems.
@@ 8.1 Mathematical definition
To formally define structural stability, we introduce the concept of topologically equivalence of dynamical systems. Two dynamical systems are called topologically equivalent if there is a homeomorphism $h:\mathbb{R}^{S}\to\mathbb{R}^{S}$ mapping their phase portraits, preserving the direction of time. Consider two smooth continuous-time dynamical systems (i) $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t))$ and (ii) $\dot{\textbf{x}}(t)=\textbf{g}(\textbf{x}(t))$. Both systems (i) and (ii) are defined in a closed region $D\in\mathbb{R}^{S}$, see Fig. 12a. System (i) is called structurally stable in a region $D_{0}\subset D$ if for any system (ii) that is sufficiently $C^{1}$-close to system (i) there are regions $U$, $V\subset D$, and $D_{0}\subset U$, $D_{0}\subset V$ such that system (i) is topologically equivalent in $U$ to system (ii) in $V$, see Fig. 12a. Here, the systems (i) and (ii) are $C^{1}$-close if their “distance,” defined as
@@
Andronov and Pontryagin offered sufficient and necessary conditions for a two-dimensional continuous-time dynamical system to be structurally stable [17]. A smooth dynamical system $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t))$ with $\textbf{x}(t)\in\mathbb{R}^{2}$, is structurally stable in a region $D_{0}\subset\mathbb{R}^{2}$ if and only if (i) it has a finite number of equilibrium points and limit cycles in $D_{0}$, and all of them are hyperbolic; (ii) there are no saddle separatrices returning to the same saddle, see Fig. 12b and c, or connecting two different saddles in $D_{0}$, see Fig. 12d. This is often called the Andronov-Pontryagin criterion, which gives the complete description of structurally stable systems on the plane. It has been proven that a typical or generic two-dimensional system always satisfies the Andronov-Pontryagin criterion and hence is structurally stable [190]. In other words, structural stability is a generic property for planar systems. Yet, this is not true for high-dimensional systems. Later on, Morse and Smale established the sufficient conditions for an $S$-dimensional dynamical systems to be structurally stable [237, 238]. Such systems, often called Morse-Smale systems, have only a finite number of equilibrium points and limit cycles, all of which are hyperbolic and satisfy a transversaility condition on their stable and unstable invariant manifolds.
Figure 12: Structural stability. (a) Andronov’s definition of structural stability. (b-d) Phase portraits of structurally unstable planar systems. This figure was redrawn from [148].
@@ 8.2 Structural stability metrics
The notion of structural stability has been explored in diverse ecological systems. For instance, Recknagel investigated the structural stability of aquatic ecological systems by using the catastrophe theory and applied his approach as an aid in decision-making for water quality management [209]. In addition, the concept of structural stability have been heavily used in soil ecological systems [97, 192, 193, 223, 258]. Below, we survey various measures that can be used to quantify the structural stability of ecological systems.
Rohr et al. introduced a mathematical framework based on the concept of structural stability to elucidate the influence of network architecture on community persistence with the GLV model [213]. They proposed that an ecological system becomes more structurally stable as the area of the parameter space of the model expands, resulting in both a dynamically stable and feasible equilibrium. In particular, the authors investigated the range of conditions necessary for the stable coexistence of all species in mutualistic systems and showed that numerous observed mutualistic network architectures tend to maximize the volume of parameter space under which species coexist. This implies that having both a nested network architecture and a small mutualistic trade-off is one of the most preferable structures for community persistence.
Grilli et al. developed a geometrical framework to study the range of conditions necessary for feasible coexistence of large ecological systems [110]. They quantified the structural stability of an ecological system using the GLV model as the volume of the feasibility region when varying intrinsic growth rates, which can be approximated by
@@
where $C_{A}$ is the connectance of the interaction matrix A, $\mu_{A}$ is the mean of the off-diagonal elements of A, and $\mu_{d}$ is the mean of the diagonal elements of A. The authors further analytically predicted the range of coexistence conditions in more than 100 empirical ecological systems. Recently, using the above approximation of structural stability, Portillo et al. explored the correlation between structural stability and various network measures, such as centrality and modularity, and illustrated that optimal modularity has a negative impact on biological diversity (structural stability) in empirical ecological systems [205].
During the same time, Saavedra et al. proposed novel metrics analogous to stabilizing niche differences and fitness differences with the GLV model, which can measure the range of conditions compatible with multi-species coexistence, i.e., structural stability [218]. The structural analogs of the niche and fitness differences are defined as
@@
Saavedra and his coauthors have utilized similar approaches to study the structural stability of various ecological systems [60, 61, 214, 219, 220, 221, 240, 241, 242]. For example, they quantified structural stability using the quantity
@@
which measures how big the deviations are from the structural vector (geometric centroid) compatible with a positive stable equilibrium [220]. They further proved that the smaller the level of global competition, the broader the conditions for having feasible solutions. In addition, Song and Saavedra proposed a measure of structural stability using the quantity $\Omega(\textbf{A})^{1/S}$ with a transformation $2\textbf{A}^{\top}\textbf{A}=\boldsymbol{\Sigma}^{-1}$, which offers an approximation to the level of external perturbations tolerated by an ecological system [242]. The authors further found that their measure is the only consistent predictor of changes in species richness in empirical ecological systems among different ecological and environmental variables.
Lately, Pettersson et al. developed a metric called instability to quantify the proximity to collapse and level of structural stability with the GLV model, which provides deep insights into the dynamics and limits of stability and collapse of ecological systems [196]. The instability metric is defined as
@@ 9 Higher-order stability analysis
Various complex networks, such as ecological networks, social networks, biological networks, and chemical reaction networks, exhibit higher-order interactions, in which the interactions go beyond pairwise relationships and involve multiple nodes or entities at the same time [14, 40, 52, 54, 64, 63, 83, 99, 264]. In particular, in ecological networks, species interactions frequently occur in higher-order combinations, where the relationship between two species is influenced by one or more additional species [27, 141, 152], see Fig. 13. From the modeling perspective, higher-order interactions can be explicitly modeled by adding higher-order terms to the classical GLV model. Alternatively, they can be implicitly captured through consumer-resource models, i.e., consumer species associated with the same resource form a higher-order interaction. While the significance of higher-order interactions has been recognized, their comprehensive impact on the stability of ecological systems has not been fully understood.
Pioneering research has shown that one type of third-order interaction, in which one species mitigates the negative interactions between two others, can stabilize well-mixed ecological systems with particular network configurations [137]. In addition, increasing the order of interactions among species can weaken the destabilizing effect and amplify the variance of species abundances at the equilibrium [78, 276]. In this section, we first introduce the GLV model with higher-order interactions and present various results related to the stability of ecological systems with higher-order interactions. We subsequently outline two potential approaches for analytically determining the stability of the GLV model with higher-order interactions via polynomial systems theory. Finally, we discuss various consumer-resource model for implicit higher-order interactions and their stability properties. Note that some results for higher-order stability also involve linear stability analysis.
@@ 9.1 GLV model with higher-order interactions
The dynamics of ecological systems with higher-order interactions are often described by the GLV model with higher-order interactions [6, 234] defined as
@@
for $i=1,2,\dots,S$. Here, $\textsf{B}\in\mathbb{R}^{S\times S\times S}$ is a third-order tensor whose off-diagonal elements represent the effect that species $j$ and $k$ has upon species $i$. $\textsf{C}\in\mathbb{R}^{S\times S\times S\times S}$ is a fourth-order tensor whose off-diagonal elements represent the effect that species $j$, $k$, $l$ has upon species $i$. In fact, the GLV model with higher-order interactions (30) belongs to the family of polynomial dynamical systems [62]. AlAdwani and Saavedra [6] determined the number of non-trivial equilibrium points of the GLV model with higher-order interactions (30) based on Bernshtein’s theorem [48] from algebraic geometry. However, the stability of those non-trivial equilibrium points remains unknown. So far, most of the stability analysis of higher-order interactions depends on linearization or numerical simulations of the GLV model with higher-order interactions or other similar models.
Figure 13: Examples of pairwise, third-order, and fourth-order interactions in ecological systems. For pairwise interactions, species 1 could produce an antibiotic, inhibiting the growth of species 2. For third-order interactions, species 3 might degrade the antibiotics produced by species 1, thus alleviating the inhibitory effect of species 1 on species 2. For fourth-order interactions, the activity of the antibiotic-degrading enzyme by species 3 may in turn be inhibited by compounds produced by species 4. This figure was redrawn from [27].
Bairey et al. found that the classical relationship between diversity and stability is reversed when considering high-order interactions by simulating the dynamics of ecological systems using a replicator model [27], which is defined based on the higher-order GLV model, with random interactions. Denote the bracket part in the GLV model with higher-order interactions (30) as $f_{i}(\textbf{x})$. The replicator model is defined as
@@
Later on, Grilli et al. explored the role of higher-order interactions in competitive ecological systems [111]. Particularly, the authors simulated different competitive ecological systems with only third-order interactions of the form:
@@
for $i=1,2,\dots,S$, which can be viewed as a special case of the higher-order GLV model. The interaction tensor B is defined from pairwise interactions and can be computed as $\textsf{B}_{ijk}=2\textbf{H}_{ij}\textbf{H}_{ik}-\textbf{H}_{ji}\textbf{H}_{jk}-\textbf{H}_{ki}\textbf{H}_{kj}$ where the first term represents the probability of species $i$ beats both species $j$ and $k$, and the remaining two terms represent the probabilities that either species $j$ or $k$ dominate. Here, the matrix H encodes the dominance relationships among species, i.e., $\textbf{H}_{ij}$ represents the probability that species $i$ outcompetes species $j$. Compared to the pairwise dynamics, the equilibrium point of (33) is unchanged, and more importantly, it is globally stable. Therefore, they concluded that incorporating higher-order interactions into competitive ecological systems enhances stability, making species coexistence resilient to perturbations of both population abundance and parameter values [111].
Singh and Baruah derived general rules for species coexistence modulated by higher-order interactions by simulating the higher-order GLV model [234]. They demonstrated that negative higher-order interactions can promote coexistence if these interactions strengthen intraspecific competition more than interspecific competition, while positive higher-order interactions can stabilize coexistence across a wide range of fitness differences, disregarding differences in strength of interspecific and intraspecific competition. Recently, Gibbs et al. also exploited the higher-order GLV model and found that when interspecific higher-order interactions became excessively harmful compared to self-regulation, the stability of coexistence in diverse ecological systems will be compromised [101]. Furthermore, they found that coexistence can also be lost when these interactions are weak and mutualistic higher-order effects became dominant [101].
@@ 9.2 Two potential approaches
As mentioned, the GLV model with higher-order interactions (30) belongs to the family of polynomial dynamical systems. Thus, polynomial systems theory can be leveraged to analytically study the stability properties of the model. We summarize two recent approaches that can analytically determine the stability of homogeneous polynomial dynamical systems [3, 62]. For simplicity, let’s assume that a higher-order ecological system with $S$ species only contains $p$th-order interactions. The resulting GLV model with higher-order interactions thus becomes homogeneous (of degree $p$), similar to the form (33). In fact, every polynomial dynamical system can be homogenized by introducing additional variables. Therefore, the two approaches can be in principle applied to analyze the stability of the GLV model with higher-order interactions.
@@ 9.2.1 Lyapunov approach
Ali and Khadir proved that the existence of a rational Lyapunov function, defined as the ratio of two polynomials, is both necessary and sufficient for asymptotic stability of a homogeneous polynomial dynamical system at the origin [3]. The rational Lyapunov function takes the form of
@@
where $r$ is a non-negative integer and $q(\textbf{x})$ is a homogeneous positive definite polynomial of degree $22r+$ 2. Since the Lyapunov inequalities on both the rational function and its derivative have sum of squares certificates, $V(\textbf{x})$ can always be found by semi-definite programming [259].

### 07 Feb 2024 v2

@@ 9 Structural stability analysis
The notion of structural stability of dynamical systems was first introduced by Andronov and Pontryagin under the name coarse (or rough) systems Andronov-1937 . Different from the previous notions of stability which consider perturbations of initial conditions for a fixed dynamical system, structural stability concerns whether the qualitative behavior of the system trajectories will be affected by small perturbations of the system model itself bavzant2000structural ; peixoto1959structural ; simitses2006fundamentals ; thom2018structural ; ziegler2013principles . In this section, we first discuss the mathematical definition of structural stability in dynamical systems. Then, we present various metrics that can be used to quantify the structural stability of ecological systems.
@@ 9.1 Mathematical definition
To formally define structural stability, we introduce the concept of topologically equivalence of dynamical systems. Two dynamical systems are called topologically equivalent if there is a homeomorphism $h:\mathbb{R}^{S}\to\mathbb{R}^{S}$ mapping their phase portraits, preserving the direction of time. Consider two smooth continuous-time dynamical systems (i) $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t))$ and (ii) $\dot{\textbf{x}}(t)=\textbf{g}(\textbf{x}(t))$. Both systems (i) and (ii) are defined in a closed region $D\in\mathbb{R}^{S}$, see Figure 13a. System (i) is called structurally stable in a region $D_{0}\subset D$ if for any system (ii) that is sufficiently $C^{1}$-close to system (i) there are regions $U$, $V\subset D$, and $D_{0}\subset U$, $D_{0}\subset V$ such that system (i) is topologically equivalent in $U$ to system (ii) in $V$, see Figure 13a. Here, the systems (i) and (ii) are $C^{1}$-close if their “distance,” defined as
@@
Andronov and Pontryagin offered sufficient and necessary conditions for a two-dimensional continuous-time dynamical system to be structurally stable Andronov-1937 . A smooth dynamical system $\dot{\textbf{x}}(t)=\textbf{f}(\textbf{x}(t))$ with $\textbf{x}(t)\in\mathbb{R}^{2}$, is structurally stable in a region $D_{0}\subset\mathbb{R}^{2}$ if and only if (i) it has a finite number of equilibrium points and limit cycles in $D_{0}$, and all of them are hyperbolic; (ii) there are no saddle separatrices returning to the same saddle, see Figure 13b and c, or connecting two different saddles in $D_{0}$, see Figure 13d. This is often called the Andronov-Pontryagin criterion, which gives the complete description of structurally stable systems on the plane. It has been proven that a typical or generic two-dimensional system always satisfies the Andronov-Pontryagin criterion and hence is structurally stable Peixoto-1962 . In other words, structural stability is a generic property for planar systems. Yet, this is not true for high-dimensional systems. Later on, Morse and Smale established the sufficient conditions for an $S$-dimensional dynamical systems to be structurally stable Smale-1961 ; Smale-1967 . Such systems, often called Morse-Smale systems, have only a finite number of equilibrium points and limit cycles, all of which are hyperbolic and satisfy a transversaility condition on their stable and unstable invariant manifolds.
Figure 13: Structural stability. (a) Andronov’s definition of structural stability. (b-d) Phase portraits of structurally unstable planar systems. This figure was redrawn from Kuznetsov-Book-04 .
@@ 9.2 Structural stability metrics
The notion of structural stability has been explored in diverse ecological systems. For instance, Recknagel investigated the structural stability of aquatic ecological systems by using the catastrophe theory and applied his approach as an aid in decision-making for water quality management recknagel1985analysis . In addition, the concept of structural stability have been heavily used in soil ecological systems garcia2012soil ; perfect1990factors ; perfect1990rates ; sankaran1999determinants ; van2007type . Below, we survey various measures that can be used to quantify the structural stability of ecological systems.
Rohr et al. introduced a mathematical framework based on the concept of structural stability to elucidate the influence of network architecture on community persistence with the GLV model rohr2014structural . They proposed that an ecological system becomes more structurally stable as the area of the parameter space of the model expands, resulting in both a dynamically stable and feasible equilibrium. In particular, the authors investigated the range of conditions necessary for the stable coexistence of all species in mutualistic systems and showed that numerous observed mutualistic network architectures tend to maximize the volume of parameter space under which species coexist. This implies that having both a nested network architecture and a small mutualistic trade-off is one of the most preferable structures for community persistence.
Grilli et al. developed a geometrical framework to study the range of conditions necessary for feasible coexistence of large ecological systems grilli2017feasibility . They quantified the structural stability of an ecological system using the GLV model as the volume of the feasibility region when varying intrinsic growth rates, which can be approximated by
@@
where $C_{A}$ is the connectance of the interaction matrix A, $\mu_{A}$ is the mean of the off-diagonal elements of A, and $\mu_{d}$ is the mean of the diagonal elements of A. The authors further analytically predicted the range of coexistence conditions in more than 100 empirical ecological systems. Recently, using the above approximation of structural stability, Portillo et al. explored the correlation between structural stability and various network measures, such as centrality and modularity, and illustrated that optimal modularity has a negative impact on biological diversity (structural stability) in empirical ecological systems portillo2022global .
During the same time, Saavedra et al. proposed novel metrics analogous to stabilizing niche differences and fitness differences with the GLV model, which can measure the range of conditions compatible with multi-species coexistence, i.e., structural stability saavedra2017structural . The structural analogs of the niche and fitness differences are defined as
@@
Saavedra and his coauthors have utilized similar approaches to study the structural stability of various ecological systems cenci2018structural ; cenci2018rethinking ; rohr2016persist ; saavedra2016seasonal ; saavedra2014structurally ; saavedra2016nested ; song2017some ; song2018guideline ; song2018structural . For example, they quantified structural stability using the quantity
@@
which measures how big the deviations are from the structural vector (geometric centroid) compatible with a positive stable equilibrium saavedra2014structurally . They further proved that the smaller the level of global competition, the broader the conditions for having feasible solutions. In addition, Song and Saavedra proposed a measure of structural stability using the quantity $\Omega(\textbf{A})^{1/S}$ with a transformation $2\textbf{A}^{\top}\textbf{A}=\boldsymbol{\Sigma}^{-1}$, which offers an approximation to the level of external perturbations tolerated by an ecological system song2018structural . The authors further found that their measure is the only consistent predictor of changes in species richness in empirical ecological systems among different ecological and environmental variables.
Lately, Pettersson et al. developed a metric called instability to quantify the proximity to collapse and level of structural stability with the GLV model, which provides deep insights into the dynamics and limits of stability and collapse of ecological systems pettersson2020predicting . The instability metric is defined as
@@ 10 Higher-order stability analysis
Various complex networks, such as ecological networks, social networks, biological networks, and chemical reaction networks, exhibit higher-order interactions, in which the interactions go beyond pairwise relationships and involve multiple nodes or entities at the same time alvarez2021evolutionary ; battiston2021physics ; billick1994higher ; boccaletti2023structure ; chen2023survey ; chen2023teasing ; dotson2022deciphering ; ghazanfar2020investigating ; wilson1992complex . In particular, in ecological networks, species interactions frequently occur in higher-order combinations, where the relationship between two species is influenced by one or more additional species bairey2016high ; kleinhesselink2022detecting ; levine2017beyond , see Figure 14. From the modeling perspective, higher-order interactions can be explicitly modeled by adding higher-order terms to the classical GLV model. Alternatively, they can be implicitly captured through consumer-resource models, i.e., consumer species associated with the same resource form a higher-order interaction. While the significance of higher-order interactions has been recognized, their comprehensive impact on the stability of ecological systems has not been fully understood.
Pioneering research has shown that one type of third-order interaction, in which one species mitigates the negative interactions between two others, can stabilize well-mixed ecological systems with particular network configurations kelsic2015counteraction . In addition, increasing the order of interactions among species can weaken the destabilizing effect and amplify the variance of species abundances at the equilibrium de2000random ; yoshino2008rank . In this section, we first introduce the GLV model with higher-order interactions and present various results related to the stability of ecological systems with higher-order interactions. We subsequently outline two potential approaches for analytically determining the stability of the GLV model with higher-order interactions via polynomial systems theory. Finally, we discuss various consumer-resource model for implicit higher-order interactions and their stability properties. Note that some results for higher-order stability also involve linear stability analysis.
@@ 10.1 GLV model with higher-order interactions
The dynamics of ecological systems with higher-order interactions are often described by the GLV model with higher-order interactions aladwani2019addition ; singh2021higher defined as
@@
for $i=1,2,\dots,S$. Here, $\textsf{B}\in\mathbb{R}^{S\times S\times S}$ is a third-order tensor whose off-diagonal elements represent the effect that species $j$ and $k$ has upon species $i$. $\textsf{C}\in\mathbb{R}^{S\times S\times S\times S}$ is a fourth-order tensor whose off-diagonal elements represent the effect that species $j$, $k$, $l$ has upon species $i$. In fact, the GLV model with higher-order interactions (32) belongs to the family of polynomial dynamical systems chen2022explicit . AlAdwani and Saavedra aladwani2019addition determined the number of non-trivial equilibrium points of the GLV model with higher-order interactions (32) based on Bernshtein’s theorem bernshtein1975number from algebraic geometry. However, the stability of those non-trivial equilibrium points remains unknown. So far, most of the stability analysis of higher-order interactions depends on linearization or numerical simulations of the GLV model with higher-order interactions or other similar models.
Figure 14: Examples of pairwise, third-order, and fourth-order interactions in ecological systems. For pairwise interactions, species 1 could produce an antibiotic, inhibiting the growth of species 2. For third-order interactions, species 3 might degrade the antibiotics produced by species 1, thus alleviating the inhibitory effect of species 1 on species 2. For fourth-order interactions, the activity of the antibiotic-degrading enzyme by species 3 may in turn be inhibited by compounds produced by species 4. This figure was redrawn from bairey2016high .
Bairey et al. found that the classical relationship between diversity and stability is reversed when considering high-order interactions by simulating the dynamics of ecological systems using a replicator model bairey2016high , which is defined based on the higher-order GLV model, with random interactions. Denote the bracket part in the GLV model with higher-order interactions (32) as $f_{i}(\textbf{x})$. The replicator model is defined as
@@
Later on, Grilli et al. explored the role of higher-order interactions in competitive ecological systems grilli2017higher . Particularly, the authors simulated different competitive ecological systems with only third-order interactions of the form:
@@
for $i=1,2,\dots,S$, which can be viewed as a special case of the higher-order GLV model. The interaction tensor B is defined from pairwise interactions and can be computed as $\textsf{B}_{ijk}=2\textbf{H}_{ij}\textbf{H}_{ik}-\textbf{H}_{ji}\textbf{H}_{jk}-\textbf{H}_{ki}\textbf{H}_{kj}$ where the first term represents the probability of species $i$ beats both species $j$ and $k$, and the remaining two terms represent the probabilities that either species $j$ or $k$ dominate. Here, the matrix H encodes the dominance relationships among species, i.e., $\textbf{H}_{ij}$ represents the probability that species $i$ outcompetes species $j$. Compared to the pairwise dynamics, the equilibrium point of (35) is unchanged, and more importantly, it is globally stable. Therefore, they concluded that incorporating higher-order interactions into competitive ecological systems enhances stability, making species coexistence resilient to perturbations of both population abundance and parameter values grilli2017higher .
Singh and Baruah derived general rules for species coexistence modulated by higher-order interactions by simulating the higher-order GLV model singh2021higher . They demonstrated that negative higher-order interactions can promote coexistence if these interactions strengthen intraspecific competition more than interspecific competition, while positive higher-order interactions can stabilize coexistence across a wide range of fitness differences, disregarding differences in strength of interspecific and intraspecific competition. Recently, Gibbs et al. also exploited the higher-order GLV model and found that when interspecific higher-order interactions became excessively harmful compared to self-regulation, the stability of coexistence in diverse ecological systems will be compromised gibbs2022coexistence . Furthermore, they found that coexistence can also be lost when these interactions are weak and mutualistic higher-order effects became dominant gibbs2022coexistence .
@@ 10.2 Two potential approaches
As mentioned, the GLV model with higher-order interactions (32) belongs to the family of polynomial dynamical systems. Thus, polynomial systems theory can be leveraged to analytically study the stability properties of the model. We summarize two recent approaches that can analytically determine the stability of homogeneous polynomial dynamical systems ahmadi2019algebraic ; chen2022explicit . For simplicity, let’s assume that a higher-order ecological system with $S$ species only contains $p$th-order interactions. The resulting GLV model with higher-order interactions thus becomes homogeneous (of degree $p$), similar to the form (35). In fact, every polynomial dynamical system can be homogenized by introducing additional variables. Therefore, the two approaches can be in principle applied to analyze the stability of the GLV model with higher-order interactions.
@@ 10.2.1 Lyapunov approach
Ali and Khadir proved that the existence of a rational Lyapunov function, defined as the ratio of two polynomials, is both necessary and sufficient for asymptotic stability of a homogeneous polynomial dynamical system at the origin ahmadi2019algebraic . The rational Lyapunov function takes the form of
@@
where $r$ is a non-negative integer and $q(\textbf{x})$ is a homogeneous positive definite polynomial of degree $22r+$ 2. Since the Lyapunov inequalities on both the rational function and its derivative have sum of squares certificates, $V(\textbf{x})$ can always be found by semi-definite programming vandenberghe1996semidefinite .